OpenStudy (anonymous):

y=2sinx, then what would be the derivative for this?

hartnn (hartnn):

whats derivative of sin x ?

OpenStudy (anonymous):

cosx

hartnn (hartnn):

2 can be pulled out of differentiation, d/dx (2 sin x) = 2 d/dx (sin x) = 2 cos x thats it!

OpenStudy (anonymous):

so we just leave the 2 alone...? so then what are you allowed to pull out of a differentiation? definitely not a variable

hartnn (hartnn):

a constant 2 is a constant

hartnn (hartnn):

\((d/dx) a f(x)= a (d/dx)f(x)=af'(x)\)

OpenStudy (anonymous):

so technically you could factor a constant out of a function being differentiated then take the derivative of the factored function, then multiply the constant back in?

hartnn (hartnn):

thats correct.

hartnn (hartnn):

but if it was something like 2+sin x then derivative of constant = 0 so, that would be 0+cos x = cos x

OpenStudy (anonymous):

oh I c ok because you have to take the derivatives seperately~ differentiation and derivative is the same thing?

hartnn (hartnn):

yes, its same thing.

OpenStudy (anonymous):

ok so what aboutfinding the derivative of sinx^2???

hartnn (hartnn):

\(\sin (x^2)\) right ? chain rule...

OpenStudy (anonymous):

OH ic~ ok thanx!

hartnn (hartnn):

what u got for sin (x^2) just to verify

OpenStudy (anonymous):

2xcosx^2?

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

oh, and how to deal with secant?

hartnn (hartnn):

d/dx (sec x) = sec x tan x

OpenStudy (anonymous):

how...?

hartnn (hartnn):

you want me to prove that ? write sec x = 1/cos x d/dx (1/cos x) =... ? use chain rule

OpenStudy (anonymous):

chain??? you mean 1(-sinx)+cosx(0)... wait that doesn't work...? hmm how to use chain rule for this...? (cosx)^-1, -cosx?????

hartnn (hartnn):

whats derivative of 1/x ?

OpenStudy (anonymous):

1(1)+x(o) 1?

hartnn (hartnn):

was that product rule you just applied ? :P

hartnn (hartnn):

derivative of x^{-1} is .... ? use the formula for derivative of x^n

OpenStudy (anonymous):

oops LOL uh, [x(0) + 1(1)]/x^2 1/x^2

hartnn (hartnn):

oh, so you tried quotient rule now....there's - *minus* in middle

hartnn (hartnn):

i would have used that d/dx (x^n) = n x^{n-1} d/dx (x^{-1}) = -1 x^{-1-1} = -1/x^2

OpenStudy (anonymous):

omg I must be tired

OpenStudy (anonymous):

oh, ok, Ic

hartnn (hartnn):

so, d/dx (1/cos x) = -1/ (cos^2 x) * (d/dx (cos x)) [using chain rule] = -1/ (cos^2 x) * (-sin x) = sin x / cos^2 x = (sin x/ cos x) * (1/ cos x) = tan x * sec x ask if any doubts....

OpenStudy (anonymous):

how did u get d/dx(cosx) in that first step? and the original question was like...d^2/d^2x just so u kno I dont understand what that notation means etiher...LOL

hartnn (hartnn):

d/dx (cos x) means derivative of cos x which is sin x

hartnn (hartnn):

d^2/d^2x = d/dx (d/dx ) means double derivative. taking derivative twice, one after another

hartnn (hartnn):

like d^2/d^2x (x^3) = d/dx (d/dx x^3) = d/dx (3x^2 ) = 6x

OpenStudy (anonymous):

oh ok so then I understand how you got the first derivative now. You saw it as (cos)^-1, which allowed for the chain rule ic. so then after tanx times secx, now it is the problem of geeting the second derivative.

hartnn (hartnn):

means you had to find d^2/d^2x (sec x) ??

OpenStudy (anonymous):

si, that would be correct. So then I tried it, and I got -secx

hartnn (hartnn):

so, now you need d/dx (sec x tan x) used product rule ? (how come the answer is so compact ? i don't think thats correct...)

OpenStudy (anonymous):

my answer is probs wrong... is not one of the answer choices on here...

OpenStudy (anonymous):

(sinx/cosx)(sinx/cosx times 1/cosx) + {1/cosx[(cosx times cosx)-(sinx times (-sinx))]}/cos^2x sin^2x/cos^3x+(cos^2x+sin^2x)/cos^3x) sin^2x+1/cos^3x -cos^2x/cos^3x -1/cosx -secx

hartnn (hartnn):

upto here its correct sin^2x+1/cos^3x

hartnn (hartnn):

how 1+sin^2 x = -cos^2 x ?

OpenStudy (anonymous):

oh!!! I think I was thinking it was sin^2x -1 LOL so then where do I go from there?

hartnn (hartnn):

you can leave it like this :(sin^2x+1)/cos^3x if there's such an option... or ((sin^2 x/ cos^2x) * (1/ cos x))+ 1/ cos^3 x = tan^2 x sec x + sec^3 x

hartnn (hartnn):

if you still didn't get the required answer, list the choices.

OpenStudy (anonymous):

got it~! oh, and I have another question, but I'm goihng to open up another question cya there?

hartnn (hartnn):

sure :)